OpenStudy (anonymous):

Which point is the vertex of the parabola y = x2+4x+1? A. (-2, -3) B. (2, 3) C. (-2, 3) D. (2, -3)

OpenStudy (anonymous):

hero :)

hero (hero):

x = -b/2a = -4/2(1) = -2 y = (-2)^2+4(-2)+1 = 4-8 +1 = 5-8 = -3 (x,y) = (-2,-3)